### v1

@@ Abstract
Transformer-based networks applied to image patches have achieved cutting-edge performance in many vision tasks. However, lacking the built-in bias of convolutional neural networks (CNN) for local image statistics, they require large datasets and modifications to capture relationships between patches, especially in segmentation tasks. Images in the frequency domain might be more suitable for the attention mechanism, as local features are represented globally. By transforming images into the frequency domain, local features are represented globally. Due to MRI data acquisition properties, these images are particularly suitable. This work investigates how the image domain (spatial or k-space) affects segmentation results of deep learning (DL) models, focusing on attention-based networks and other non-convolutional models based on MLPs. We also examine the necessity of additional positional encoding for Transformer-based networks when input images are in the frequency domain. For evaluation, we pose a skull stripping task and a brain tissue segmentation task. The attention-based models used are PerceiverIO and a vanilla Transformer encoder. To compare with non-attention-based models, an MLP and ResMLP are also trained and tested. Results are compared with the Swin-Unet, the state-of-the-art medical image segmentation model. Experimental results show that using k-space for the input domain can significantly improve segmentation results. Also, additional positional encoding does not seem beneficial for attention-based networks if the input is in the frequency domain. Although none of the models matched the Swin-Unet’s performance, the less complex models showed promising improvements with a different domain choice.
@@ 3.2 Implementation Details
This study uses Python (version 3.10.12), PyTorch (version 2.0.1), TorchIO (version 0.18.92), Ray (version 2.5.1), and PyTorch Lightning (version 2.0.5) for all implementations and analyses  [^1] The source code is publicly available at https://www.github.com/rauschecker-sugrue-labs/kspace-segmentation. To propagate the slices of each sample through the network in the frequency domain, we employ the 2D real FFT implementation in PyTorch. This method operates under the assumption that the outcome of the FFT is Hermitian symmetric, which holds true in our case since we are using already reconstructed MRI data. By exploiting this symmetry, the input is halved. As the real part and imaginary part of the resulting complex numbers are saved separately in two vectors, the size of the input remains the same even after the Fourier transformation. The training routine for all non-baseline models is designed to allow independent specification of the target domain for inputs and labels at startup. This flexibility enables the definition of various combinations of input and label domains. When labels are in the spatial domain, the task becomes a classification problem and the model is trained using a cross-entropy loss; when labels are transformed into the frequency domain, it becomes a regression problem and the model is trained using mean squared error loss.
@@ 3.3 Segmentation Performance in different Domains
Table 1: Performance metrics in spatial domain, k-space and k-space-to-spatial domain for different models on skull stripping.
Domain/Metric Spatial K-space K-Space $\rightarrow$ Spatial
Dice Recall Spec. Dice Recall Spec. Dice Recall Spec.
@@
Transformer encoder 0.971 0.971 0.992 0.971 0.969 0.993 0.972 0.973 0.993
nnU-Net 0.986 0.987 0.996 - - - - - -
Swin-Unet 0.994 0.995 0.974 - - - - - -
@@
0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Swin-UnetnnU-NetResMLPMLPTransformerPerceiverIO0.920.920.920.920.960.960.960.960.880.880.880.880.630.630.630.630.70.70.70.70.420.420.420.420.820.820.820.820.690.690.690.690.790.790.790.790.460.460.460.460.880.880.880.880.680.680.680.680.860.860.860.860.420.420.420.42Dice scoreSpatialK-SpaceK-Space $\rightarrow$ Spatial
Figure 1: Quantitative comparison of all models on brain tissue segmentation in varies domains.
@@
Table 2: Attention-based models with and without positional encoding on skull stripping and brain tissue segmentation in k-space. Average Dice score is reported for brain tissue segmentation task.
Model Metric Skull Stripping Brain Tissue Segmentation
@@
PerceiverIO Dice 0.927 0.930 0.461 0.463
Transformer Dice 0.971 0.970 0.790 0.790
@@ 3.4 Complexity comparison with State-of-the-Art Models
The training and testing of the nnU-Net and Swin-Unet in spatial domain shows that both models outperform the other models in any domain constellation. However, in skull stripping the differences are relatively small. The reason for this is most likely because this task is not complex enough to see differences between domains. Looking at the results of the brain tissue segmentation in Fig. 1, the differences become more obvious. In this analysis, it becomes clear that nnU-Net and the Swin-Unet exhibits the most effective performance in brain tissue segmentation. However, both the Transformer encoder in the k-space-to-spatial domain and ResMLP offer competitive segmentation outcomes. This becomes particularly clear when comparing the complexity of these models. Table 3 shows the floating point operations (FLOPs) required for the forward and backward pass of the models used. The number of model parameters is also documented. The results clearly show that although the Swin-Unet outperforms the other non-baseline models, it is also significantly more complex. Comparing the Transformer encoder with the Swin-Unet, it can be seen that more than twice as many FLOPs are required for the forward and backward pass. In addition, the ResMLP requires less then 600×$\times{}10^{9}$ FLOPs for both passes. Taking this into consideration, the performance of the Transformer encoder and ResMLP is impressive. In summary, the choice of domain significantly impacts the results for brain tissue segmentation. The correct domain can compensate for a lack of model capacity and thus good segmentation results can be achieved with simple models.
@@ 4 Conclusion and Limitations
Some aspects of this work limit the expressiveness and applicability of the results obtained. One aspect is that only the binary cross-entropy loss, cross-entropy loss and mean squared error loss were considered as loss functions in this work. Other loss functions such as Dice Loss or Focal Loss would also be worth considering. The OASIS dataset is freely accessible and offers isotropic resolution, which simplifies our pre-processing pipeline. However, it should also be noted that this dataset contains presumably healthy subjects, which makes the segmentation task easier. Also, the approach used in this work cannot be directly translated into the clinical setting. By using the real 2D FFT it was assumed that the input is Hermitian symmetric. However, this assumption cannot be made for raw MRI k-space data.

### v3

@@ Abstract
Transformer-based networks applied to image patches have achieved cutting-edge performance in many vision tasks. However, lacking the built-in bias of convolutional neural networks (CNN) for local image statistics, they require large datasets and modifications to capture relationships between patches, especially in segmentation tasks. Images in the frequency domain might be more suitable for the attention mechanism, as local features are represented globally. By transforming images into the frequency domain, local features are represented globally. Due to MRI data acquisition properties, these images are particularly suitable. This work investigates how the image domain (spatial or k-space) affects segmentation results of deep learning (DL) models, focusing on attention-based networks and other non-convolutional models based on MLPs. We also examine the necessity of additional positional encoding for Transformer-based networks when input images are in the frequency domain. For evaluation, we pose a skull stripping task and a brain tissue segmentation task. The attention-based models used are PerceiverIO and a vanilla Transformer encoder. To compare with non-attention-based models, an MLP and ResMLP are also trained and tested. Results are compared with the Swin-Unet, the state-of-the-art medical image segmentation model. Experimental results indicate that using k-space for the input domain can significantly improve segmentation results. Also, additional positional encoding does not seem beneficial for attention-based networks if the input is in the frequency domain. Although none of the models matched the Swin-Unet’s performance, the less complex models showed promising improvements with a different domain choice.
@@ 3.2 Implementation Details
This study uses Python (version 3.10.12), PyTorch (version 2.0.1), TorchIO (version 0.18.92), Ray (version 2.5.1), and PyTorch Lightning (version 2.0.5) for all implementations and analyses  [^1] The source code is publicly available at https://www.github.com/rauschecker-sugrue-labs/kspace-segmentation. To propagate the slices of each sample through the network in the frequency domain, we employ the 2D real FFT implementation in PyTorch. This method operates under the assumption that the outcome of the FFT is Hermitian symmetric, which holds true in our case since we are using already reconstructed MRI data. By exploiting this symmetry, the input is halved. As the real part and imaginary part of the resulting complex numbers are saved separately in two vectors, the size of the input remains the same even after the Fourier transformation. The training routine for all non-baseline models is designed to allow independent specification of the target domain for inputs and labels at startup. This flexibility enables the definition of various combinations of input and label domains. When labels are in the spatial domain, the task becomes a classification problem and the model is trained using a cross-entropy loss; when labels are transformed into the frequency domain by using the 2D FFT, it becomes a regression problem and the model is trained using mean squared error loss. We use three domain configurations. In the spatial domain, both input and output are in the spatial domain. In the k-space domain, the input is transformed into the frequency domain using FFT, and the output is predicted in the frequency domain. In the k-space-to-spatial domain, the input is in the frequency domain, and the output is predicted in the spatial domain.
@@ 3.3 Segmentation Performance in different Domains
Table 1: Dice similarity coefficient (DSC), sensitivity (sens) and specificity (spec) for different models on skull-stripping in spatial domain, k-space and k-space-to-spatial domain.
Spatial K-Space K-Space to Spatial
DSC Sens Spec DSC Sens Spec DSC Sens Spec
@@
Trans. Encoder 0.971 0.971 0.992 0.971 0.969 0.993 0.972 0.973 0.993
nnU-Net 0.986 0.987 0.996
Swin-UNet 0.994 0.995 0.974
@@
0.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Swin-UnetnnU-NetResMLPMLPTrans. EncoderPerceiverIODice scoreSpatialK-SpaceK-Space $\rightarrow$ Spatial
Figure 1: Mean and standard deviations of the Dice score among all models on brain tissue segmentation in spatial, k-space, and k-space-to-spatial domain. The mean values are represented by the bars and the standard deviations are indicated by the error bars.
@@
Table 2: Dice scores for attention-based models with and without positional encoding (PE) on skull stripping and brain tissue segmentation in k-space. Average Dice score is reported for brain tissue segmentation task.
Architecture Skull Stripping Brain Tissue Segmentation
@@
PerceiverIO 0.927 0.930 0.461 0.463
Trans. Encoder 0.971 0.970 0.790 0.790
@@ 3.4 Complexity comparison with State-of-the-Art Models
The training and testing of the nnU-Net and Swin-Unet in spatial domain shows that both models outperform the other models in any domain constellation. However, in skull stripping the differences are relatively small. The reason for this is most likely because this task is not complex enough to see differences between domains. Looking at the results of the brain tissue segmentation in Fig. 1, the differences become more obvious. In this analysis, it becomes clear that nnU-Net and the Swin-Unet exhibits the most effective performance in brain tissue segmentation. However, both the Transformer encoder in the k-space-to-spatial domain and ResMLP offer competitive segmentation outcomes. This becomes particularly clear when comparing the complexity of these models. Table 3 shows the floating point operations (FLOPs) required for the forward and backward pass of the models used. The number of model parameters is also documented. The results clearly show that although the Swin-Unet outperforms the other non-baseline models, it is also much more complex. Comparing the Transformer encoder with the Swin-Unet, it can be seen that more than twice as many FLOPs are required for the forward and backward pass. In addition, the ResMLP requires less then 600×$\times{}10^{9}$ FLOPs for both passes. Taking this into consideration, the performance of the Transformer encoder and ResMLP is impressive. In summary, the choice of domain significantly impacts the results for brain tissue segmentation. The correct domain can compensate for a lack of model capacity and thus good segmentation results can be achieved with simple models.
@@ 4 Conclusion and Limitations
Some aspects of this work limit the scope and applicability of the results obtained. One aspect is that only the binary cross-entropy loss, cross-entropy loss and mean squared error loss were considered as loss functions in this work. Other loss functions such as Dice Loss or Focal Loss would also be worth considering. The OASIS dataset is freely accessible and offers isotropic resolution, which simplifies our pre-processing pipeline. However, it should also be noted that this dataset contains presumably healthy subjects, which makes the segmentation task easier. Also, the approach used in this work cannot be directly translated into the clinical setting. By using the real 2D FFT it was assumed that the input is Hermitian symmetric. However, this assumption cannot be made for raw MRI k-space data.